Fresh / Soph / Jr / Sr / Grad (circle one) College and Major:

The grade I expect to receive is:

36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 2 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 3 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Mid-

4 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Semester

5 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Eval-

6 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

u-

7 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

a-

8 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tion

9 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 10 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 11 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation Sam-12 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation pling, 13 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation Sur-14 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation veys 15 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation & 16 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation So-17 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 18 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation ety 19 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation February 20 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 21 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 22 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

• 23 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

This 24 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

eval-25 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

u-26 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

a-27 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tion 28 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

form 29 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

has 30 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

only 31 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

one 32 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

pur-33 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

pose:

34 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

to 35 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

im-36 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

prove 37 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 38 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

qual-39 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ity 40 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

of 41 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

in-42 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

struc-43 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tion 44 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

of 45 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

this 46 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

course, 47 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

both 48 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

for 49 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 50 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

rest 51 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

of 52 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

this 53 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

semester, 54 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

and 55 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

for 56 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

fu-57 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ture 58 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

semesters.

59 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

• 60 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Below 61 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

are 62 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

sev-63 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

eral 64 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

gen-65 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

eral 66 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

state-67 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ments 68 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

about 69 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 70 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

course.

71 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Please 72 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

cir-73 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

cle 74 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

your 75 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

re-76 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

sponse 77 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

to 78 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

each 79 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

state-80 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ment 81 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

us-82 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ing 83 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 84 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 85 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

point 86 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

scale 87 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

in-88 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 89 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

cated.

90 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

• 91 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Use 92 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 93 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

blank 94 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

space 95 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

on 96 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 97 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

next 98 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

page, 99 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

to 100 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

of-101 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

fer 102 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

fur-103 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ther 104 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

com-105 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ments 106 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

& 107 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

sug-108 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ges-109 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tions 110 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

on 111 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 112 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

course.

113 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

• 114 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Your 115 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

frank 116 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

com-117 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ments 118 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

and 119 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

con-120 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

struc-121 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tive 122 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

crit-123 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 124 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

cism 125 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

would 126 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

be 127 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

greatly 128 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ap-129 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

pre-130 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 131 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ated.

132 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 133 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Do 134 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

not 135 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

sign 136 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

your 137 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

name.

138 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation Overall 139 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation My 140 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation un-141 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation der-142 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation stand-143 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation ing 144 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation of 145 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation the 146 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation ma-147 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation te-148 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation rial 149 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation cov-150 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation ered 151 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation to 152 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation date:

153 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 154

36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

1  2  3  4  5

I’m totally lost Challenging, but I think I’m understanding the material This course is really easy 155 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation My 156 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation in-157 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation ter-158 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation est 159 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation in, 160 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation and 161 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation en-162 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation joy-163 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation ment 164 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation of, 165 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation this 166 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation course 167 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation to 168 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation date:

169 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 170

1  2  3  4  5

I’m bored stiff It’s OK but I like another course better One of my favorite classes 171 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation Course 172 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 173 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

The course is not too hard and not too easy.

174

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

175 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

If 176 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

you 177 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

dis-178 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

agree, 179 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

is 180 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

it 181 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

too 182 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

\underline{hard} 183 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

or 184 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

too 185 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

\underline{easy}?

186 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 187 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 188 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 189 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

The course is not too slow and not too fast.

190

1 2 3 4 5

191 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

If 192 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

you 193 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

dis-194 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

agree, 195 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

is 196 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

it 197 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

too 198 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

\underline{slow} 199 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

or 200 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

too 201 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

\underline{fast}?

202 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 203 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 204 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 205 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

The things I am learning in this course will be useful in other classes or outside of school.

206

1 2 3 4 5

207 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 208 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 209 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 210 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

Please 211 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

use 212 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

this 213 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

page 214 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

to 215 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

pro-216 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

vide 217 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

any 218 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

other 219 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

com-220 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ments 221 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

or 222 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

sug-223 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ges-224 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tions 225 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

on 226 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 227 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

course 228 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

(pos-229 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 230 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tive 231 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

or 232 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

neg-233 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

a-234 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tive)

235 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

that 236 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

you 237 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

think 238 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

would 239 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

be 240 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

use-241 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

ful.

242 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

You 243 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

might 244 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

think 245 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

about 246 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 247 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

in-248 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

struc-249 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tor, 250 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 251 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

TA, 252 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 253 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

lec-254 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

tures, 255 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 256 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

text-257 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

books, 258 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 259 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

home-260 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

works, 261 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 262 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

group 263 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

projects, 264 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 265 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

class-266 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

room 267 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

setup, 268 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

or 269 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

any 270 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

other 271 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

as-272 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

pect 273 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

of 274 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

the 275 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation

course.

276 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 277 36-303: Sampling, Surveys & SocietyMid-Semester Evaluation 278